# Internal Dynamics of Multiple Populations in 28 Galactic Globular Clusters: A Wide-Field study with Gaia and the Hubble Space Telescope

### v1

@@ Abstract
We present the first comprehensive analysis of the internal dynamics of multiple stellar populations (MPs) in 28 Galactic Globular Clusters (GCs) across a wide field of view, extending from the innermost regions to the clusters’ outskirts. Using astro-photometric catalogs from ground-based observations, Gaia, and the Hubble Space Telescope (HST), we identify first- (1P) and second-population (2P) stars, and study the internal dynamics of MPs using high-precision Gaia DR3 and HST proper motions. Our results reveal that while the 1P transitions from isotropy to slight tangential anisotropy toward the outer regions, 2P stars become increasingly radially anisotropic beyond the half-light radius. We also explore the connection between the dynamics of MPs and the clusters’ structural and dynamical properties, finding statistically significant differences in the anisotropy profiles of dynamically young and non-relaxed clusters, particularly beyond the 1-2 half-light radii. In these regions, 1P stars transition from isotropic to slightly tangentially anisotropic motion, while 2P stars become more radially anisotropic. In contrast, dynamically older clusters, with mixed MPs, exhibit weaker relative differences. Furthermore, clusters with orbits closer to the Galactic center exhibit larger dynamical differences between 1P and 2P stars than those with larger peri-Galactic radii. These findings are consistent with a scenario where 2P stars form in a more centrally concentrated environment, where the interaction with the Milky Way tidal field plays a crucial role in the dynamical evolution of MPs, especially of 1P.
@@
More than two decades of Hubble Space Telescope (HST) observations have revealed that nearly all Galactic Globular clusters (GCs) host Multiple Populations (MPs) with at least two main groups of stars: first-population (1P) and second-population (2P Gratton et al., 2012; Piotto et al., 2015; Renzini et al., 2015; Milone et al., 2017; Bastian & Lardo, 2018). These groups exhibit specific chemical compositions, with 2P stars being for example enriched in Sodium, Nitrogen and Helium and depleted in Carbon and Oxygen (see e.g. Bastian & Lardo, 2018; Milone & Marino, 2022, for recent reviews). Despite several models have been proposed to explain the formation of MPs, none can simultaneous match the numerous observational constraints (see e.g. Renzini et al., 2015; Bastian & Lardo, 2018; Milone & Marino, 2022). Some scenarios predict that 2P stars formed from the ejecta of short-lived massive 1P stars (Ventura et al., 2001; Decressin et al., 2007; D’Ercole et al., 2010; Denissenkov & Hartwick, 2014). As the expelled gas accumulated toward the cluster center, 2P stars would dominate the composition of the central cluster regions. Alternatively, the chemical composition of 2G stars could be the result of the accretion of chemically enriched material emitted from 1P stars in binary configuration (de Mink et al., 2009; Bastian et al., 2013; Gieles et al., 2018).
@@
Over the past decade, a number of studies have analyzed the spatial distribution and the internal kinematics of MPs in various GCs, using HST photometry and proper motions (see e.g. Libralato et al., 2023; Leitinger et al., 2023), MUSE/APOGEE line-of-sight (LoS) velocities (Szigeti et al., 2021; Martens et al., 2023), ground-based photometry coupled with Gaia proper motions, or a combination of these methods (Milone et al., 2018; Cordoni et al., 2020a, b, 2023; Cadelano et al., 2024). Such works presented the first evidence of significant dynamical differences between 1P and 2P stars in some GCs, qualitatively aligning with -body simulations’ results. However, the limited number of clusters and the small field of view prevented a complete and detailed characterization of the phenomenon.
In this work we aim to overcome the limitation of previous studies by investigating the internal dynamics of MPs in a sample of 28 GCs, combining photometric information from HST (Nardiello et al., 2018), ground-based (see e.g. Stetson et al., 2019; Jang et al., 2022) and Gaia (Mehta et al., 2024) and accurate HST (Libralato et al., 2022) and Gaia (Gaia Collaboration et al., 2023a) proper motions. This work presents the first comprehensive dynamical investigation of MPs in a large sample of clusters, extending from the central arcminutes to nearly the tidal radius. The paper is structured as follows: in Sec. 2 we introduce the dataset and MPs selection process. Section 3 presents the study of the internal dynamics in the analyzed clusters and the global dynamical profiles of MPs. Finally, the discussion and summary are presented in Sec. 4 and 5, respectively.
@@ 2 Data
The final selection of 1P and 2P stars for each cluster and the analyzed field of view are shown in App. A as online supplementary material. 333The sample of analyzed clusters also includes type-II GCs, i.e. clusters with internal variation of heavy-elements (see e.g. Milone et al., 2017; Marino et al., 2019). However, considering the small number of stars belonging to the anomalous populations, i.e. enriched stars, we only study stars with light-elements abundance variations.
@@ 2.1 Multiple Populations from ground-based photometry
To select 1P and 2P stars from the ChMs we adapted the procedure introduced in Milone et al. (2017) on HST ChMs to our ground-based ChMs, as done in Jang et al., in preparation. Figure 1 illustrates the procedure for the cluster NGC 2808. Briefly, we first selected by eye the bulk of 1P stars among the stars with , and fitted a straight line to 1P stars (black solid line in Fig. 1, leftmost panel). We then rotated the ChMs by the angle defined by slope of the 1P best-fit line. We refer to the new rotated reference frame as and (Fig. 1, middle panel). By construction, 1P stars define a horizontal distribution in vs , and we can use the distribution to disentangle 1P and 2P stars. We used the scikit-learn python package (Pedregosa et al., 2011) and employed Gaussian Mixture Models (GMM) to fit 2 Gaussians to the distribution. We remind here that GMM is applied directly to the data points, so that the results do not depend on the bin choice. Finally, 1P stars (red points in the rightmost panel) have been selected as stars with , with being the intersection between the two best-fit Gaussians, indicated by the horizontal line in the middle panel of Fig. 1. 2P stars are marked with blue points. Additionally, stars with unusual values of and have been excluded from the selection.
@@ 2.2 Multiple populations from HST photometry
To study MPs among RGB stars in the innermost cluster regions we exploited the ChMs and selection carried out in Milone et al. (2017). In a nutshell, HST ChMs have been derived by means of the verticalized vs. (for the abscissa) and vs. 444 CMDs (on the ordinate). 1P and 2P stars have been selected with the procedures detailed in Milone et al. (2017, see e.g. Sec. 3.3 for a detailed description). Each HST ChM has then been cross matched with the proper motion catalogs of Libralato et al. (2022).
@@ 3 Internal dynamics of Multiple Populations
After carefully selecting MPs in the 28 analyzed clusters, we exploited HST (Libralato et al., 2022) and Gaia DR3 (Gaia Collaboration et al., 2023a) proper motions to investigate the internal dynamics of stars belonging to different populations. Specifically, we followed the procedure described in Libralato et al. (2022, 2023, for HST) and Bianchini et al. (2018)555see also Milone et al. (2018); Bianchini et al. (2019); Vasiliev (2019); Cordoni et al. (2020a, b); Vasiliev & Baumgardt (2021); Cordoni et al. (2023); Cadelano et al. (2024).. We first transformed the celestial coordinates and their proper motion components 666 into a Cartesian reference frame , using the orthographic projection (eq. 2 in Gaia Collaboration et al., 2018), and then the proper motions into their sky-projected radial and tangential components, defined as and , with pointing outward (positive) or inward (negative), and positive when counterclockwise. and here indicate the proper motion of each star relative to the motion of the cluster, determined by Vasiliev & Baumgardt (2021). Furthermore, considering that the sample of analyzed stars is different from that analyzed in Vasiliev & Baumgardt (2021), we repeated the analysis determining the mean cluster motion from our sample of stars, and accounting for systematic errors as in Vasiliev & Baumgardt (2021)777routines available at https://github.com/GalacticDynamics-Oxford/GaiaTools., finding consistent results.
@@ 3.1 Individual dynamical profiles
In this section, we present the study of the dynamical profiles of 1P and 2P stars in a few individual clusters with sufficiently large samples, namely NGC 0104, NGC 2808, NGC 5904, and NGC 6205. The dynamical profile of the remaining clusters are shown in App. B as online supplementary material.
Figure 2 shows, from top to bottom, the mean radial and tangential profiles, the radial and tangential dispersion, and the anisotropy profiles, defined as . The red and blue solid lines indicate the dynamical quantities, determined as described in Sec. 3, while the shaded rectangles indicate the relative uncertainties and the extension of each radial bin. The distance from the cluster center are in units of indicated in the middle panel (Harris, 1996, revision of 2010). Black diamonds indicate the values of the 1-dimensional dispersion from Libralato et al. (2022), while the regions with available HST data is indicated by the vertical green lines. The top and right axis (shown in gray) of each panel display distances in units of parsec and km/s, converted considering cluster distances determined in Baumgardt & Hilker (2018).
@@
Concerning the dispersion profiles, 1P stars exhibit a larger tangential and similar radial dispersion compared with 2P stars, especially outside , while 1P and 2P have similar radial dispersion profiles. As a result, 1P stars exhibit an isotropic motion in the central regions, possibly shifting toward tangentially anisotropic motion in the outer regions. On the other hand, the motion of 2P stars is consistent with being isotropic in the cluster center, becoming radially anisotropic beyond 1. A similar pattern is also observed in NGC 3201, consistently with Cadelano et al. (2024). The mean, dispersion and anisotropy profiles of the remaining clusters are displayed in App. B. Overall, we find a good agreement between HST and Gaia dynamical profiles in overlapping or neighbouring regions.
@@ 3.2 Global dynamical profiles
Given the low number of 1P and 2P stars in many clusters, a detailed analysis of the individual dynamical profiles as a function of cluster radius is often unfeasible. Therefore, similarly to Libralato et al. (2023), we opted to investigate the internal dynamics in all clusters simultaneously, combining together the dispersion profiles of the individual clusters888We find worth mentioning that Libralato et al. (2023) normalized positions and proper motions of cluster stars and then derived the dispersion profiles. However, due to the non-relative nature of Gaia proper motions, we cannot use the same procedure on our dataset. Hence, we first derive the dynamical profiles in each cluster, and then normalize and combine the profiles.. In order to properly compare different clusters, we normalized the radial coordinates to the cluster from Harris (1996, revision of 2010) and the dispersion profiles to the central dispersion determined in Libralato et al. (2022) from HST data. Four clusters do not have determination of the central velocity dispersion from HST data, namely NGC 1904, NGC 4147, NGC 6712 and NGC 7492. For these clusters, we adopted the central dispersion determined in Vasiliev & Baumgardt (2021). We remind here that such approach can be used only to compare dispersion profiles, but not mean motions, as the latter are not on a relative scale.
@@
The normalized global dispersion and anisotropy profiles as a function of distance from cluster center for the 28 analyzed GCs are shown in Fig. 3. Global profiles for 1P and 2P stars are shown in the first and second columns, while the comparison between the average trend is shown in the third column. The average profiles were determined using the Locally Weighted Regression (LOESS) as implemented in Cappellari et al. (2013). To estimate the uncertainties in the average trend, we repeated the LOESS fitting on 1000 bootstrapped samples, drawing the values of in each realization from a Gaussian distribution centered on the observed value with a dispersion equal to the observed uncertainties. The uncertainties were determined as the 16th and 84th percentiles of the 1000 LOESS fits. This is a more robust approach than simply bootstrapping the sample, as it accounts for observed uncertainties as well. Finally, we estimated the significance of the observed differences between 1P and 2P stars accounting for uncertainties in both data and fitting. The details of the procedure are outlined in App. C. In a nutshell, for each of the 1000 realizations of the LOESS fit, we computed the difference between 1P and 2P as a function of the radial coordinate. The 1, 2 and 3 confidence regions are shown as gray shaded regions in the rightmost panels, where we display the difference between 1P and 2P a function of normalized radius. Additionally, we quantified the statistical significance of each point as the fraction of simulations which returned a difference smaller than the observed one, at the same radial location. The color of the line in the rightmost panels is indicative of of the statistical significance, in units of , as indicated in the right colorbar.
@@
While there are no differences between the global radial dispersion profiles of 1P and 2P stars, 2P stars exhibit a lower tangential dispersion with respect to 1P stars, especially in the outer regions, e.g. . As a consequence, the 1P displays an isotropic/tangentially anisotropic motion in the outer regions, while the 2P is radially anisotropic. Due to the large uncertainties of the mean trend, the observed differences are significant beyond the 3 level only between 2 and 4 . Nonetheless, the magnitude of the differences is consistent with the predictions from the theoretical simulations (see e.g. Tiongco et al., 2019; Vesperini et al., 2021). Additionally, the global profiles within are consistent with the results of Libralato et al. (2023).

### v2

@@ Abstract
We present a detailed analysis of the internal dynamics of multiple stellar populations (MPs) in 28 Galactic Globular Clusters (GCs) across a wide field of view, extending from the innermost regions to the clusters’ outskirts. Using astro-photometric catalogs from ground-based observations, Gaia, and the Hubble Space Telescope (HST), we identify first- (1P) and second-population (2P) stars, and study the internal dynamics of MPs using high-precision Gaia DR3 and HST proper motions. Our results reveal that while the 1P transitions from isotropy to slight tangential anisotropy toward the outer regions, 2P stars become increasingly radially anisotropic beyond the half-light radius. We also explore the connection between the dynamics of MPs and the clusters’ structural and dynamical properties, finding statistically significant differences in the anisotropy profiles of dynamically young and non-relaxed clusters, particularly beyond the 1-2 half-light radii. In these regions, 1P stars transition from isotropic to slightly tangentially anisotropic motion, while 2P stars become more radially anisotropic. In contrast, dynamically older clusters, with mixed MPs, exhibit weaker relative differences. Furthermore, clusters with orbits closer to the Galactic center exhibit larger dynamical differences between 1P and 2P stars than those with larger peri-Galactic radii. These findings are consistent with a scenario where 2P stars form in a more centrally concentrated environment, where the interaction with the Milky Way tidal field plays a crucial role in the dynamical evolution of MPs, especially of 1P.
@@
More than two decades of Hubble Space Telescope (HST) observations have revealed that nearly all Galactic Globular clusters (GCs) host Multiple Populations (MPs) with at least two main groups of stars: first-population (1P) and second-population (2P Gratton et al., 2012; Piotto et al., 2015; Renzini et al., 2015; Milone et al., 2017; Bastian & Lardo, 2018). These groups exhibit specific chemical compositions, with 2P stars being for example enriched in Sodium, Nitrogen and Helium and depleted in Carbon and Oxygen (see e.g. Bastian & Lardo, 2018; Milone & Marino, 2022, for recent reviews). Despite several models have been proposed to explain the formation of MPs, none can simultaneous match the numerous observational constraints (see e.g. Renzini et al., 2015; Bastian & Lardo, 2018; Milone & Marino, 2022). Some scenarios predict that 2P stars formed from the ejecta of short-lived massive 1P stars (Ventura et al., 2001; Decressin et al., 2007; D’Ercole et al., 2010; Denissenkov & Hartwick, 2014). As the expelled gas accumulated toward the cluster center, 2P stars would dominate the composition of the central cluster regions. Alternatively, the chemical composition of 2G stars could be the result of the accretion of chemically enriched material emitted from 1P stars in binary configuration (de Mink et al., 2009; Bastian et al., 2013) or supermassive stars (Gieles et al., 2018).
@@
Over the past decade, a number of studies have analyzed the spatial distribution and the internal kinematics of MPs in various GCs, using HST photometry and proper motions (see e.g. Libralato et al., 2023; Leitinger et al., 2023), MUSE/APOGEE line-of-sight (LoS) velocities (Szigeti et al., 2021; Martens et al., 2023), ground-based photometry coupled with Gaia proper motions, or a combination of these methods (Milone et al., 2018; Cordoni et al., 2020a, b, 2023; Cadelano et al., 2024). Such works presented the first evidence of significant dynamical differences between 1P and 2P stars in some GCs, qualitatively aligning with -body simulations’ results. However, the limited number of clusters and the small field of view prevented a complete and detailed characterization of the phenomenon. More recently, Dalessandro et al. (2024) and Leitinger et al. (2023) analyzed the 3D kinematics of 16 and 30 GCs, respectively. Both studies combined ground-based photometric data with Gaia proper motions and LoS velocities to investigate the rotation and dynamics of MPs. Dalessandro et al. (2024) found differences in the rotation of 1P and 2P stars, with the latter exhibiting stronger rotation. In contrast, Leitinger et al. (2023) detected significant rotational differences only in NGC 0104 (47 Tuc). Dalessandro et al. (2024) also applied analytical models to examine internal dynamics, finding differences in the anisotropy between first-generation and second-generation stars.
In this work we aim to overcome the limitation of previous studies by investigating the internal dynamics of MPs in a sample of 28 GCs, combining photometric information from HST (Nardiello et al., 2018), ground-based (see e.g. Stetson et al., 2019; Jang et al., 2022) and Gaia (Mehta et al., 2024) and accurate HST (Libralato et al., 2022) and Gaia (Gaia Collaboration et al., 2023a) proper motions. By doing this we improve the identification of 1P and 2P stars by means of the ChMs, and extend the analyzed field of view from the innermost cluster regions to their outskirt. The paper is structured as follows: in Sec. 2 we introduce the dataset and MPs selection process. Section 3 presents the study of the internal dynamics in the analyzed clusters and the global dynamical profiles of MPs. Finally, the discussion and summary are presented in Sec. 4 and 5, respectively.
@@ 2 Data
The final selection of 1P and 2P stars for each cluster and the analyzed field of view are shown in App. A as online supplementary material.
@@ 2.1 Multiple Populations from ground-based photometry
To select 1P and 2P stars from the ChMs we adapted the procedure introduced in Milone et al. (2017) on HST ChMs to our ground-based ChMs, as done in Jang et al., in review. Figure 1 illustrates the procedure for the cluster NGC 2808. Briefly, we first selected by eye the bulk of 1P stars among the stars with , and fitted a straight line to 1P stars (black solid line in Fig. 1, leftmost panel). We then rotated the ChMs by the angle defined by slope of the 1P best-fit line. We refer to the new rotated reference frame as and (Fig. 1, middle panel). By construction, 1P stars define a horizontal distribution in vs , and we can use the distribution to disentangle 1P and 2P stars. We used the scikit-learn python package (Pedregosa et al., 2011) and employed Gaussian Mixture Models (GMM) to fit 2 Gaussians to the distribution. We remind here that GMM is applied directly to the data points, so that the results do not depend on the bin choice. Finally, 1P stars (red points in the rightmost panel) have been selected as stars with , with being the intersection between the two best-fit Gaussians, indicated by the horizontal line in the middle panel of Fig. 1. 2P stars are marked with blue points. Additionally, stars with unusual values of and have been excluded from the selection. On average, less than 1% of stars have been excluded, and the results are dynamical results are unaffected.
@@ 2.2 Multiple populations from HST photometry
To study MPs among RGB stars in the innermost cluster regions we exploited the ChMs and selection carried out in Milone et al. (2017). In a nutshell, HST ChMs have been derived by means of the verticalized vs. (for the abscissa) and vs. 333 CMDs (on the ordinate). 1P and 2P stars have been selected with the procedures detailed in Milone et al. (2017, see e.g. Sec. 3.3 for a detailed description). Each HST ChM has then been cross matched with the proper motion catalogs of Libralato et al. (2022).
@@ 2.3 Multiple populations from Gaia XP synthetic photometry
2.4 Anomalous stars Type II clusters
Recent studies using HST photometry and high-resolution spectroscopy have shown that about 20% of Galactic GCs exhibit variations in [Fe/H], s-process elements, and total C+N+O (e.g., Marino et al., 2015, 2019; Yong et al., 2015; Milone et al., 2017; McKenzie et al., 2022). These clusters are commonly referred to as Type II , with their heavy-element-enriched stars labeled as “anomalous”. Photometrically, these clusters show additional sequences in their ChMs, running parallel to the “canonical” 1P/2P stars.
Four of the clusters analyzed in this study, namely NGC 1261, NGC 1851, NGC 6934, and NGC 7089, are classified as Type II clusters by Milone et al. (2017) and Marino et al. (2019). To identify anomalous stars, we used the methodology outlined in Milone et al. (2017) on both HST and ground-based ChMs. These stars are highlighted with orange points in Figures 5–7, and the number of anomalous stars is indicated in Tab. 1. For NGC 6934, however, we could only identify anomalous stars using the HST ChM.
@@ 3 Internal dynamics of Multiple Populations
After carefully selecting MPs in the 28 analyzed clusters, we exploited HST (Libralato et al., 2022) and Gaia DR3 (Gaia Collaboration et al., 2023a) proper motions to investigate the internal dynamics of stars belonging to different populations. Specifically, we followed the procedure described in Libralato et al. (2022, 2023, for HST) and Bianchini et al. (2018)444see also Milone et al. (2018); Bianchini et al. (2019); Vasiliev (2019); Cordoni et al. (2020a, b); Vasiliev & Baumgardt (2021); Cordoni et al. (2023); Cadelano et al. (2024).. We first transformed the celestial coordinates and their proper motion components 555 into a Cartesian reference frame , using the orthographic projection (eq. 2 in Gaia Collaboration et al., 2018), and then the proper motions into their sky-projected radial and tangential components, defined as and , with pointing outward (positive) or inward (negative), and positive when counterclockwise. and here indicate the proper motion of each star relative to the motion of the cluster, determined by Vasiliev & Baumgardt (2021). Furthermore, considering that the sample of analyzed stars is different from that analyzed in Vasiliev & Baumgardt (2021), we repeated the analysis determining the mean cluster motion from our sample of stars, and accounting for systematic errors as in Vasiliev & Baumgardt (2021)666routines available at https://github.com/GalacticDynamics-Oxford/GaiaTools., finding consistent results.
@@ 3.1 Individual dynamical profiles
In this section, we present the study of the dynamical profiles of 1P and 2P stars in a few individual clusters with sufficiently large samples, namely NGC 0104, NGC 2808, NGC 5904, and NGC 6205. The dynamical profile of the remaining clusters are shown in App. B as online supplementary material. Results for anomalous stars, if present, are shown in orange.
Figure 2 shows, from top to bottom, the mean radial and tangential profiles, the radial and tangential dispersion, and the anisotropy profiles, defined as . Red and blue colours indicate the results for 1P and 2P, determined as described in Sec. 3, while the shaded rectangles indicate the relative uncertainties and the extension of each radial bin. The distance from the cluster center are in units of indicated in the middle panel (Harris, 1996, revision of 2010). Black diamonds indicate the values of the 1-dimensional dispersion from Libralato et al. (2022), while the regions with available HST data is indicated by the vertical green lines. The top and right axis (shown in gray) of each panel display distances in units of parsec and km/s, converted considering cluster distances determined in Baumgardt & Hilker (2018).
@@
Concerning the dispersion profiles, 1P stars exhibit a larger tangential and similar radial dispersion compared with 2P stars, especially outside . As a result, 1P stars exhibit an isotropic motion in the central regions, possibly shifting toward tangentially anisotropic motion in the outer regions. On the other hand, the motion of 2P stars is consistent with being isotropic in the cluster center, becoming radially anisotropic beyond 1. A similar pattern is also observed in NGC 3201, consistently with Cadelano et al. (2024). The mean, dispersion and anisotropy profiles of the remaining clusters are displayed in App. B, with anomalous stars in Type II GCs, namely NGC 1261, NGC 1851, NGC 6934 and NGC 7089, shown with orange markers777It is worth noting that the dynamical quantities of anomalous stars are determined from a limited sample of stars in all clusters except for NGC 1851.. Overall, we find a good agreement between HST and Gaia dynamical profiles in overlapping or neighboring regions.
@@ 3.2 Global dynamical profiles
Given the low number of 1P and 2P stars in many clusters, a detailed analysis of the individual dynamical profiles as a function of cluster radius is often unfeasible. Therefore, similarly to Libralato et al. (2023), we opted to investigate the internal dynamics combining together all clusters. To this goal, we used two different methods: 1) creating stacked 1P/2P catalogs by combining all stars from different clusters, and 2) combining the dispersion profiles of individual clusters to compute global profiles. For the former, the sample was divided into equally populated radial bins, each containing at least 200 stars, to derive global profiles in the same way as individual clusters.
@@
In order to properly compare different clusters, we normalized the radial coordinates to the cluster from Harris (1996, revision of 2010) and the dispersion profiles to the central dispersion determined in Libralato et al. (2022) from HST data. Four clusters do not have determination of the central velocity dispersion from HST data, namely NGC 1904, NGC 4147, NGC 6712 and NGC 7492. For these clusters, we adopted the central dispersion determined in Vasiliev & Baumgardt (2021). We remind here that such approach can be used only to compare dispersion profiles, but not mean motions, as the latter are not on a relative scale. Additionally, combining individual stars from all clusters, we subtracted each cluster’s overall tangential motion profile (derived in Vasiliev & Baumgardt, 2021) to account for differences in the rotational pattern of each cluster888For NGC 7492, whose internal dynamical profiles were not studied in Vasiliev & Baumgardt (2021), we subtracted the cluster’s mean tangential motion..
In the following, we show the results obtained with approach 1), i.e. combining individual cluster stars, for 1P and 2P. The profiles derived from stacked 1P/2P profiles, along with supplementary materials and figures, are publicly available on the GitHub and as a Streamlit application. We also include tests, described in the following, on the influence of individual clusters on the global profiles.
@@
The normalized global dispersion and anisotropy profiles as a function of distance from cluster center for the 28 analyzed GCs are shown in Fig. 3. Global profiles for 1P and 2P stars are shown in the first and second columns, while the comparison between the average trend is shown in the third column. The global profiles were determined using the Locally Weighted Regression (LOESS) as implemented in Cappellari et al. (2013). To estimate the uncertainties in the average trend, we repeated the LOESS fitting on 1000 bootstrapped samples, drawing the values of in each realization from a Gaussian distribution centered on the observed value with a dispersion equal to the observed uncertainties. The uncertainties were determined as the 16th and 84th percentiles of the 1000 LOESS fits. This is a more robust approach than simply bootstrapping the sample, as it accounts for observed uncertainties as well. Finally, we estimated the significance of the observed differences between 1P and 2P stars accounting for uncertainties in both data and fitting. The details of the procedure are outlined in App. C. In a nutshell, for each of the 1000 realizations of the LOESS fit, we computed the difference between 1P and 2P as a function of the radial coordinate. The 1, 2 and 3 confidence regions are shown as gray shaded regions in the rightmost panels, where we display the difference between 1P and 2P a function of normalized radius. Additionally, we quantified the statistical significance of each point as the fraction of simulations which returned a difference smaller than the observed one, at the same radial location. The color of the line in the rightmost panels is indicative of of the statistical significance, in units of , as indicated in the right colorbar.
@@
While there are no differences between the global radial dispersion profiles of 1P and 2P stars, 2P stars exhibit a lower tangential dispersion compared to 1P stars, especially in the outer regions (e.g., ). Consequently, the 1P displays isotropic or tangentially anisotropic motion in these regions, while the 2P is radially anisotropic. The observed differences are significant beyond between 2 and . Our global profiles within are also consistent with the conclusions of Libralato et al. (2023). Additionally, although we computed the global profile of anomalous stars, % of these stars belong to NGC 1851, making the profile unrepresentative of the entire cluster sample. Therefore, we choose not to present it. Nevertheless, we note that the profile is qualitatively consistent with 2P global profiles.
To identify and investigate potential biases in the global profiles caused by certain clusters contributing more to the stacked datasets (e.g., the most populated clusters), we conducted the following tests: i) derived global profiles using only the six most populated clusters (NGC 0104, NGC 2808, NGC 5272, NGC 6205, NGC 5904, and NGC 7089); ii) repeated the analysis excluding these six clusters; and iii) randomly sampled an equal number of 1P and 2P stars (50 stars, or the total number of available stars for clusters with fewer than 50 1P/2P) from each cluster to create 1,000 combined dataset realizations. In all three tests, we obtained results consistent with those shown in Fig. 3 (i.e. radially anisotropic 2P and isotropic/tangentially anisotropic 1P), with the main difference being the increased uncertainties resulting from smaller samples. Furthermore, for each stacked dataset, we generated heat maps to illustrate the contribution of each cluster in each radial bin and found no significant biases. To further validate our findings, we recomputed the global profiles from the stacked individual dynamical profiles (e.g. the profiles presented in Figs. 10-12), again obtaining consistent results. The results of these tests and their descriptions are available on GitHub and as a Streamlit application.